OpenStudy (anonymous):

What fraction of a sample of cobalt-60 (half-life = 5 years) will remain after 25 years?

OpenStudy (wolf1728):

Let's assume you start with 100 grams. (Incidentally, the half-life of cobalt 60 which is 5.2714 years (wikipedia)) But we'll stick with 5 years. According to the formula here: http://www.1728.org/halflife.htm Ending Amount = Beginning Amount / (2^n) where n= number of half lives In 25 years, cobalt 60 will undergo 5 half-lives. So, Ending Amount = 100g / 2^5 Ending Amount = 100g / 32 Ending Amount = 3.125 g So, basically, after 5 half-lives (or 25 years) the amount of cobalt 60 remaining is .03125 times the original amount.

OpenStudy (dumbcow):

another way of looking at it is : \[\large (\frac{1}{2})^{\frac{t}{5}}\] let t = 25

OpenStudy (wolf1728):

Okay, that solves it too :-)